A space station in the form of a large wheel, 185 m in diameter, rotates to provide an “artificial gravity” of 9.2 m/s^2 for people located on the outer rim. Find the rotational frequency of the wheel that will produce this effect. Answer in units of rpm.
Answer #1

here,

diameter of large wheel , d = 185 m

radius , r = d/2 = 92.5 m

let the rotational frequency of wheel be w

the artificial gravity , g = w^2 * r

g = w^2 * r

9.2 = w^2 * 92.5

solving for w

w = 0.315 Hz

the rotational frequency of wheel is 0.315 Hz